# GA-SmaAt-GNet: Generative Adversarial Small Attention GNet for Extreme Precipitation Nowcasting

### v1

@@ 3.2 Training
We conduct a comparative evaluation of our proposed GA-SmaAt-GNet and SmaAt-GNet against the SmaAt-UNet model SmaAtUnet and the persistence model. The persistence model uses the last input image of the input sequence repeated 12 times as its prediction.
Training is performed on the precipitation dataset specified in Section 4.1. The maximum epoch limit is set to 200, although in practice, this limit was never reached. To enhance training efficiency and prevent overfitting, we employed the same early stopping criterion as in SmaAtUnet , where training halts if there is no improvement on the validation set for more than 15 epochs.
For all models, we initialized the learning rate at 0.001, following the lines of SmaAtUnet . To prevent overfitting of GA-SmaAt-GNet’s discriminator, we adjusted the backpropagation frequency of the discriminator to once every two iterations. Additionally, a learning rate scheduler was implemented, reducing the learning rate to a tenth of its size when no improvement occurred on the validation set for four consecutive epochs. GA-SmaAt-GNet featured separate learning rate schedulers for the discriminator and generator, monitoring their respective losses on the validation set. The Adam optimizer is utilized as optimizer adam with a batch size of 32 and Mean Squared Error (MSE) served as the loss function for SmaAt-UNet and SmaAt-GNet. For GA-SmaAt-GNet, the loss function in Eq. (3) is used with an empirically found $\lambda$-value of $10^{6}$. All models are trained on an Apple M2 Ultra with 128GB of memory. The implementation of the models is available on GitHub [^1] https://github.com/EloyReulen/GA-SmaAt-GNet where there is also more information about pre-trained models, which are available upon request.
@@ 3.3 Evaluation
MSE serves as the metric for assessing the performance of all models in terms of regression and is calculated as follows:
@@
where $n$ represents the number of samples, $y_{i}$ is the ground truth and $\hat{y}_{i}$ is the model’s prediction. Here, $y_{i}$ and $\hat{y}_{i}$ are tensors of size $12\times 64\times 64$ and $[y_{i}-\hat{y}_{i}]^{2}_{j}$ indicates the $j$th element of the tensor $[y_{i}-\hat{y}_{i}]^{2}$. In addition to MSE, we evaluate model performance using various binary classification metrics. To this end, precipitation maps are transformed into binary representations by denormalizing and summing up 12 consecutive frames, where 12 is the number of output frames of the models. Subsequently, values exceeding specific thresholds (0.5, 10, and 20 mm/h) are set to one, while others are set to zero. This enables performance assessment for both higher and lower precipitation intensities. The thresholds of 0.5 mm/h (used to distinguish between rain and no rain) aligns with SmaAtUnet ; convLSTM . From these binary maps, true positives (TP) (prediction=1, target=1), false positives (FP) (prediction=1, target=0), true negatives (TN) (prediction=0, target=0), and false negatives (FN) (prediction=0, target=1) are calculated. Binary classification metrics used include F1 score, Critical Success Index (CSI), Heidke Skill Score (HSS), and Matthews Correlation Coefficient (MCC).
@@ 4.1 Dataset
We train and test the models on 25 years of precipitation data from the Royal Netherlands Meteorological Institute (Koninklijk Nederlands Meteorologisch Instituut, KNMI). The dataset is a climatological radar rainfall dataset of 5 minute precipitation depths at a 2.4-km grid, which have been adjusted with validated and complete rain gauge data from both KNMI rain gauge networks. The first network is an automated rain gauge network of 32 stations and the second network is a manual rain gauge network of about 325 stations. The collected dataset contains data from 1998 till 2022. Of all 25 years of data, years 1998-2016 are used for the training set and years 2017-2022 are used for the test set.
@@
The input for the models is a sequence of 12 precipitation maps which corresponds to one hour of precipitation data. This input length allows for a fair comparison with SmaAt-UNet SmaAtUnet which also used 12 precipitation maps as input. The output consists of the next 12 precipitation maps. As described in SmaAtUnet many of the precipitation maps contain little to no rain data which can cause the models to become biased towards predicting zero values. To prevent this from happening, only sequences of which the output frames contain more than 50% rainy pixels are selected similar to the approach by SmaAtUnet . This 50% is calculated only on the pixels that contain precipitation data. The sample sizes of the preprocessed dataset and the original dataset are compared in Table 1. The preprocessed dataset is available upon request, for more details see the GitHub page of this paper.
@@ 5 Results and discussion
Following the training procedure detailed in Section 3.2, the model with the lowest validation error (MSE) for each architecture was chosen for performance evaluation on the precipitation test set. For each test case, the models are run 10 times with unfixed dropout layers. The mean of these 10 results serves as the prediction. This process makes it also possible to obtain uncertainty for the models and will be discussed in more detail in Section 5.2.1. Additionally, all models are compared against the SmaAt-UNet model and the persistence model which serves as the baseline.
@@ 5.1 Precipitation nowcasting evaluation
The obtained results on the precipitation test set are presented in Table 2 and Fig. 4. It is important to note that these results are computed after denormalization, similar to those outlined by SmaAtUnet . The results in Table 2 reveal that all models perform significantly better than the persistence model in terms of Mean Squared Error (MSE). Furthermore, we observe that the introduced SmaAt-GNet architecture leads to a slight reduction in MSE on the test set when compared to SmaAt-UNet. However, the lowest MSE of all models is observed for GA-SmaAt-GNet.
In addition to the overall MSE on the test set, we computed MSE per predicted time step, as illustrated in Fig. 4. It is evident that all models exhibit similar behavior as time progresses, i.e., MSE increases for predictions further into the future. However, when comparing SmaAt-GNet to SmaAt-UNet, it becomes clear that the overall performance gain of SmaAt-GNet can primarily be attributed to the first half-hour of the prediction.
@@
For a threshold of 0.5 mm/h, SmaAt-GNet demonstrates superior performance in all binary metrics, closely followed by SmaAt-UNet and GA-SmaAt-GNet. GA-SmaAt-GNet performs similar to the SmaAt-UNet in terms of F1 and CSI and slightly worse in terms of HSS and MCC. However all models, significantly outperform the persistence model at this threshold.
At a heavier precipitation intensity with a threshold of 10 mm/h, the persistence model performs exceptionally poor for all binary metrics. SmaAt-GNet demonstrates an improvement over SmaAt-UNet for all metrics at this threshold. GA-SmaAt-GNet exhibits superior performance compared to all other models, outperforming them for all metrics.
In the scenario of extreme precipitation, defined by an accumulated rainfall exceeding 20 mm in an hour, one can observe that SmaAt-UNet performs worse than the persistence model for all metrics. GA-SmaAt-GNet demonstrates superior performance for all binary metrics. In addition, we note that SmaAt-GNet significantly outperforms the persistence model and SmaAt-UNet in terms all binary metrics, however it is not able to reach the level of performance of the GA-SmaAt-GNet.
@@
Furthermore, we compute binary metrics per season to provide insights into model performance across different seasons. Fig. 5 illustrates the binary metrics per season for a precipitation intensity of 10 mm/h. This threshold was chosen as no significant differences were observed for threshold of 0.5 mm/h, and for a threshold of 20 mm/h some seasons experienced a shortage of positive cases, making it less suitable for meaningful comparisons. We note that across all metrics, the tested models tend to perform less effectively in winter compared to other the other seasons. This can possibly be attributed to the scarcity of positive heavy precipitation cases in the training set during winter. Moreover, heavy precipitation in winter tends to have a different nature, being characterized by prolonged and widespread events whereas heavy summer precipitation is often local and short-lived kmni_extreme . When comparing SmaAt-UNet and SmaAt-GNet, we note an interesting observation. SmaAt-UNet exhibits slightly superior performance in spring and autumn for all metrics, while SmaAt-GNet notably outperforms in summer. This is particularly interesting as summer is the season marked by the most intense rain showers kmni_extreme . However, the standout performer in Fig. 5 is GA-SmaAt-GNet which significantly outperforms all models for most metrics in all seasons with the exception of winter.
Figure 7: Epistemic uncertainty quantification results per time step calculated on the precipitation test set. The mean epistemic uncertainties over all time steps are indicated by the dotted lines. The persistence model is not shown as it has no uncertainty.
@@
Refer to caption (b) 10 min ahead Refer to caption (c) 30 minutes ahead Refer to caption (d) 50 minutes ahead
@@
We also provide a visual comparison of the models on a specific case from the test set where the accumulated rainfall exceeded 20 mm. The model outputs for this case are depicted in Fig. 6. As anticipated, when compared with the ground truth, all models exhibit decreased accuracy as the predicted frame extends further into the future. Upon closer examination of the last two frames predicted by SmaAt-UNet, we note the following observation: it overestimates the rain intensity in the bottom left corner more than the other models. Additionally, in the prediction, the high intensity rain cluster spreads out over time and is almost unnoticeable in the prediction at the 60 minute mark, while in the ground truth the intensity stays the roughly same.
@@ 5.2 Uncertainty analysis
In Bayesian modeling the consideration of uncertainty is critical and it is present in two primary forms uncertanties . First, aleatoric uncertainty which is the inherent noise present in observations, such as sensor noise. This form of uncertainty persists even when the volume of data is increased. Aleatoric uncertainty is further classified into homoscedastic uncertainty, which remains constant across different inputs, and heteroscedastic uncertainty, where the degree of uncertainty depends on the inputs to the model. Heteroscedastic aleatoric uncertainty holds particular significance in computer vision applications.
@@ 5.2.1 Epistemic uncertainty
We approximate epistemic uncertainty by employing test-time dropout (TTD), a method used for estimating Bayesian inference, as demonstrated by aatransunet ; dropout_brain . With TTD, dropout layers are not fixed at test time like they normally would be. This intentional randomness during testing provides insights into the model’s uncertainty. We perform TTD for $k=10$ epochs, generating a posterior distribution which results in 10 sets of 12 precipitation maps. Subsequently, we calculate uncertainty maps by determining the variance across the 10 precipitation maps for each of the 12 predicted timesteps, resulting in 12 uncertainty maps, each corresponding to a predicted timestep. If, instead of variance, we calculate the mean, we obtain the model’s prediction. Additionally, if we take the mean of all uncertainty maps generated by the model we get an overall measure of uncertainty on the entire test set.
@@
Figure 11: Activation heatmaps obtained using Grad-CAM on all pixels predicted with rain after binarization. The one hour accumulated precipitation of the input, ground truth and model prediction are shown in (a). The activation heatmaps of the encoders can be found in (b) and the activation heatmaps of the decoder can be found in (c). Each row represents a level (depth) of the encoder or decoder of the GA-SmaAt-GNet generator. For the encoders we show activation heatmaps for the double depthwise-seperable convolution (DSC) blocks and CBAM layers and for the decoder we show the activation heatmaps of the DSC blocks.
Furthermore, we examine the epistemic uncertainty across seasons using the GA-SmaAt-GNet model, as illustrated in Fig. 9. An interesting finding comes to light, indicating that epistemic uncertainty reaches its highest point during the summer season. This coincides with the observation that mean precipitation is at its peak during this season, as depicted in Fig. 8(c). The increased occurrence of sudden, heavy rain showers during summer, as noted in kmni_extreme , is likely a contributing factor to the heightened epistemic uncertainty. Notably, it is intriguing to observe that, despite higher precipitation levels in winter, epistemic uncertainty tends to be lower compared to spring. This phenomenon may be attributed to the specific precipitation patterns characterizing winter, involving longer-lasting, lower-intensity rainfall events on a larger scale, as highlighted in kmni_extreme . Consequently, these patterns suggest that epistemic uncertainty is more influenced by precipitation intensity than the overall amount, supporting our observations derived from the illustrative examples showcased in Fig. 8.
@@ 5.2.2 Aleatoric uncertainty
To obtain heteroscedastic aleatoric uncertainty we require a neural network that maps the input to both an output and a measure of aleatoric uncertainty given by predicted variance, $\hat{\sigma}_{i}^{2}$, as described by Kendall et al. uncertanties . The predicted variance $\hat{\sigma}_{i}^{2}$ is then implicitly learned from the loss function. Instead of learning $\hat{\sigma}_{i}^{2}$ directly, the log variance $s_{i}:=\log{\hat{\sigma}_{i}^{2}}$ is learned to improve training stability by avoiding a potential division by 0. The loss function that is used to obtain aleatoric uncertainty can be defined as:

### v2

@@ 3.2 Training
We conduct a comparative evaluation of our proposed GA-SmaAt-GNet and SmaAt-GNet against SmaAt-UNet [24], RainNet [26] and the persistence model. The persistence model uses the last input image of the input sequence repeated 12 times as its prediction and RainNet is a well-established CNN that advanced precipitation nowcasting with deep learning models.
Training is performed on the precipitation dataset specified in Section 4.1. The maximum epoch limit is set to 200. To enhance training efficiency and prevent overfitting, we employed the same early stopping criterion as in [24], where training halts if there is no improvement on the validation set for more than 15 epochs.
For all models except RainNet [26], we initialized the learning rate at 0.001, following the lines of [24]. For RainNet, an initial learning rate of 0.0001 was utilized, aligning with the parameters specified in the original study. To prevent overfitting of GA-SmaAt-GNet’s discriminator, we adjusted the backpropagation frequency of the discriminator to once every two iterations. Additionally, a learning rate scheduler was implemented, reducing the learning rate to a tenth of its size when no improvement occurred on the validation set for four consecutive epochs. GA-SmaAt-GNet featured separate learning rate schedulers for the discriminator and generator, monitoring their respective losses on the validation set. The Adam optimizer is utilized as optimizer [43] with a batch size of 32 and Mean Squared Error (MSE) served as the loss function for SmaAt-UNet and SmaAt-GNet. For GA-SmaAt-GNet, the loss function in Eq. (3) is used with an empirically found $\lambda$-value of $10^{6}$. For RainNet, the logcosh function was employed as the loss function, consistent with the approach used in the original study [26]. All models are trained on an Apple M2 Ultra with 128GB of memory. The implementation of the models is available on GitHub [^1] https://github.com/EloyReulen/GA-SmaAt-GNet where there is also more information about pre-trained models, which are available upon request.
@@ 3.3 Evaluation
The Mean Squared Error (MSE) is used as the standard metric for evaluating the regression performance of all models, a common practice in meteorological regression analyses [24, 30, 34]. Additionally, its sensitivity to extreme values, due to the squared term, makes it particularly suitable for our focus on extreme precipitation. MSE is calculated as follows:
@@
where $n$ represents the number of samples, $y_{i}$ is the ground truth and $\hat{y}_{i}$ is the model’s prediction. Here, $y_{i}$ and $\hat{y}_{i}$ are tensors of size $12\times 64\times 64$ and $[y_{i}-\hat{y}_{i}]^{2}_{j}$ indicates the $j$th element of the tensor $[y_{i}-\hat{y}_{i}]^{2}$.
In addition to MSE, we evaluate model performance using various binary classification metrics. To this end, precipitation maps are transformed into binary representations by denormalizing and summing up 12 consecutive frames, where 12 is the number of output frames of the models. Subsequently, values exceeding specific thresholds (0.5, 10, and 20 mm/h) are set to one, while others are set to zero. This enables performance assessment for both higher and lower precipitation intensities. The thresholds of 0.5 mm/h (used to distinguish between rain and no rain) aligns with [24, 14]. From these binary maps, true positives (TP) (prediction=1, target=1), false positives (FP) (prediction=1, target=0), true negatives (TN) (prediction=0, target=0), and false negatives (FN) (prediction=0, target=1) are calculated. Binary classification metrics used include F1 score, Critical Success Index (CSI), Heidke Skill Score (HSS), and Matthews Correlation Coefficient (MCC).
@@ 4.1 Dataset
We train and test the models on 25 years of precipitation data from the Royal Netherlands Meteorological Institute (Koninklijk Nederlands Meteorologisch Instituut, KNMI [^2] https://www.knmi.nl/home). The dataset is a climatological radar rainfall dataset of 5 minute precipitation depths at a 2.4-km grid, which have been adjusted with validated and complete rain gauge data from both KNMI rain gauge networks. The first network is an automated rain gauge network of 32 stations and the second network is a manual rain gauge network of about 325 stations. The collected dataset contains data from 6 February 1998 8:05 till 31 December 2022 23:55. Of all 25 years of data, years 1998-2016 are used for the training set and years 2017-2022 are used for the test set.
@@
The input for the models is a sequence of 12 precipitation maps which corresponds to one hour of precipitation data. This input length allows for a fair comparison with SmaAt-UNet [24] which also used 12 precipitation maps as input. The output consists of the next 12 precipitation maps. As described in [24] many of the precipitation maps contain little to no rain data which can cause the models to become biased towards predicting zero values. To prevent this from happening, only sequences of which the output frames contain more than 50% rainy pixels are selected similar to the approach by [24]. This 50% is calculated only on the pixels that contain precipitation data. The sample sizes of the preprocessed dataset and the original dataset are compared in Table 1. The preprocessed dataset is available upon request, for more details see the GitHub page of this paper.
@@ 5 Results and discussion
Following the training procedure detailed in Section 3.2, the model with the lowest validation error (MSE) for each architecture was chosen for performance evaluation on the precipitation test set. For each test case, the models are run 10 times with unfixed dropout layers. The mean of these 10 results serves as the prediction. This process makes it also possible to obtain uncertainty for the models and will be discussed in more detail in Section 5.2.1. Additionally, all models are compared against SmaAt-UNet [24], RainNet [26] and the persistence model which serves as the baseline.
@@ 5.1 Precipitation nowcasting evaluation
The obtained results on the precipitation test set are presented in Table 2 and Fig. 4. It is important to note that these results are computed after denormalization, similar to those outlined by [24]. The results in Table 2 reveal that all models perform significantly better than the persistence model in terms of Mean Squared Error (MSE). Furthermore, we observe that the introduced SmaAt-GNet architecture leads to a slight reduction in MSE on the test set when compared to SmaAt-UNet and RainNet. However, the lowest MSE of all models is observed for GA-SmaAt-GNet.
In addition to the overall MSE on the test set, we computed MSE per predicted time step, as illustrated in Fig. 4. It is evident that all models exhibit similar behavior as time progresses, i.e., MSE increases for predictions further into the future. However, when comparing SmaAt-GNet to SmaAt-UNet, it becomes clear that the overall performance gain of SmaAt-GNet can primarily be attributed to the first half-hour of the prediction. Another notable observation is that RainNet performs quite well for the initial 15 minutes of the prediction as it is only outperformed by GA-SmaAt-GNet, but as the lead time increases it is outperformed by all other models.
@@
For a threshold of 0.5 mm/h, RainNet demonstrates superior performance in all binary metrics, closely followed by SmaAt-GNet, SmaAt-UNet and GA-SmaAt-GNet in that order. GA-SmaAt-GNet performs similar to the SmaAt-UNet in terms of F1 and CSI and slightly worse in terms of HSS and MCC. However, all models significantly outperform the persistence model at this threshold.
At a heavier precipitation intensity with a threshold of 10 mm/h, the persistence model performs exceptionally poor for all binary metrics. Among the deep learning models, RainNet performs the worst across all metrics at this threshold. Furthermore, SmaAt-GNet demonstrates an improvement over SmaAt-UNet for all metrics at this threshold. GA-SmaAt-GNet exhibits superior performance compared to all other models, outperforming them for all metrics.
In the scenario of extreme precipitation, defined by an accumulated rainfall exceeding 20 mm in an hour, one can observe that SmaAt-UNet and RainNet perform worse than the persistence model for all metrics. This may be due to high spatial smoothing, leading to underestimation of precipitation in extreme cases. This effect could arise from minimizing an objective function like Mean Squared Error (MSE) or logcosh, which tends to smooth out high-intensity values. The persistence model does not experience this issue since it simply repeats the precipitation intensity of the input. GA-SmaAt-GNet demonstrates superior performance for all binary metrics. In addition, we note that SmaAt-GNet significantly outperforms the persistence model and SmaAt-UNet in terms all binary metrics, however it is not able to reach the level of performance of the GA-SmaAt-GNet.
@@
Furthermore, we compute binary metrics per season to provide insights into model performance across different seasons. Fig. 5 illustrates the binary metrics per season for a precipitation intensity of 10 mm/h. This threshold was chosen as no significant differences were observed for threshold of 0.5 mm/h, and for a threshold of 20 mm/h some seasons experienced a shortage of positive cases, making it less suitable for meaningful comparisons. We note that across all metrics, the tested models tend to perform less effectively in winter compared to other the other seasons. This can possibly be attributed to the scarcity of positive heavy precipitation cases in the training set during winter. Moreover, heavy precipitation in winter tends to have a different nature, being characterized by prolonged and widespread events whereas heavy summer precipitation is often local and short-lived [3]. Among the deep learning models, RainNet consistently demonstrates the lowest performance across all metrics in every season. Notably, it is even outperformed by the persistence model in winter and summer. When comparing SmaAt-UNet and SmaAt-GNet, we note an interesting observation. SmaAt-UNet exhibits slightly superior performance in spring and autumn for all metrics, while SmaAt-GNet notably outperforms in summer. This is particularly interesting as summer is the season marked by the most intense rain showers [3]. However, the standout performer in Fig. 5 is GA-SmaAt-GNet which significantly outperforms all models for most metrics in all seasons with the exception of winter.
Figure 7: Epistemic uncertainty quantification results per time step calculated on the precipitation test set. The mean epistemic uncertainties over all time steps are indicated by the dotted lines. The persistence model is not shown as it has no uncertainty. RainNet is not included as it utilizes a different dropout technique than the other models.
@@
(b) 10 min ahead
(c) 30 minutes ahead
(d) 50 minutes ahead
@@
We also provide a visual comparison of the models on a specific case from the test set where the accumulated rainfall exceeded 20 mm. The model outputs for this case are depicted in Fig. 6. As anticipated, when compared with the ground truth, all models exhibit decreased accuracy as the predicted frame extends further into the future.
When analyzing the predictions made by RainNet, we notice that they exhibit more structural detail compared to those generated by other models. This enhanced detail is likely a result of the different loss function used during RainNet’s training process. Additionally, we observe that the intensity of high-precipitation area tends to decrease quite rapidly in RainNet’s prediction.
Upon closer examination of the last two frames predicted by SmaAt-UNet, we note the following observation: it overestimates the rain intensity in the bottom left corner more than the other models. Additionally, in the prediction, the high intensity rain cluster spreads out over time and is almost unnoticeable in the prediction at the 60 minute mark, while in the ground truth the intensity stays the roughly same.
@@ 5.2 Uncertainty analysis
In Bayesian modeling the consideration of uncertainty is critical and it is present in two primary forms [44]. First, aleatoric uncertainty which is the inherent noise present in observations, such as sensor noise. This form of uncertainty persists even when the volume of data is increased. Aleatoric uncertainty is further classified into homoscedastic uncertainty, which remains constant across different inputs, and heteroscedastic uncertainty, where the degree of uncertainty depends on the inputs to the model. Heteroscedastic aleatoric uncertainty holds particular significance in computer vision applications.
@@ 5.2.1 Epistemic uncertainty
We approximate epistemic uncertainty by employing test-time dropout (TTD), a method used for estimating Bayesian inference, as demonstrated by [28, 40]. With TTD, dropout layers are not fixed at test time like they normally would be. This intentional randomness during testing provides insights into the model’s uncertainty. We perform TTD for $k=10$ epochs, generating a posterior distribution which results in 10 sets of 12 precipitation maps. Subsequently, we calculate uncertainty maps by determining the variance across the 10 precipitation maps for each of the 12 predicted time steps, resulting in 12 uncertainty maps, each corresponding to a predicted time step. If, instead of variance, we calculate the mean, we obtain the model’s prediction. Additionally, if we take the mean of all uncertainty maps generated by the model we get an overall measure of uncertainty on the entire test set.
@@
Figure 11: Activation heatmaps obtained using Grad-CAM on all pixels predicted with rain after binarization. The one hour accumulated precipitation of the input, ground truth and model prediction are shown in (a). The activation heatmaps of the encoders can be found in (b) and the activation heatmaps of the decoder can be found in (c). Each row represents a level (depth) of the encoder or decoder of the GA-SmaAt-GNet generator. For the encoders we show activation heatmaps for the double depthwise-separable convolution (DSC) blocks and CBAM layers and for the decoder we show the activation heatmaps of the DSC blocks.
Furthermore, we examine the epistemic uncertainty across seasons using the GA-SmaAt-GNet model, as illustrated in Fig. 9. An interesting finding comes to light, indicating that epistemic uncertainty reaches its highest point during the summer season. This coincides with the observation that mean precipitation is at its peak during this season, as depicted in Fig. 9(c). The increased occurrence of sudden, heavy rain showers during summer, as noted in [3], is likely a contributing factor to the heightened epistemic uncertainty. Notably, it is intriguing to observe that, despite higher precipitation levels in winter, epistemic uncertainty tends to be lower compared to spring. This phenomenon may be attributed to the specific precipitation patterns characterizing winter, involving longer-lasting, lower-intensity rainfall events on a larger scale, as highlighted in [3]. Consequently, these patterns suggest that epistemic uncertainty is more influenced by precipitation intensity than the overall amount, supporting our observations derived from the illustrative examples showcased in Fig. 8.
@@ 5.2.2 Aleatoric uncertainty
To obtain heteroscedastic aleatoric uncertainty we require a neural network that maps the input to both an output and a measure of aleatoric uncertainty given by predicted variance, $\hat{\sigma}_{i}^{2}$, as described by Kendall et al. [44]. The predicted variance $\hat{\sigma}_{i}^{2}$ is then implicitly learned from the loss function. Instead of learning $\hat{\sigma}_{i}^{2}$ directly, the log variance $s_{i}:=\log{\hat{\sigma}_{i}^{2}}$ is learned to improve training stability by avoiding a potential division by 0. The loss function that is used to obtain aleatoric uncertainty can be defined as: